OpenStudy (anonymous):

If the dimensions of a rectangle become one-fourth of the original dimensions, then area of the new rectangle will be _______of the area of the original rectangle.

OpenStudy (anonymous):

Do they give you choices?

OpenStudy (anonymous):

Nope I have to type in the answer

OpenStudy (anonymous):

Oh, alright. So, what it seems like to me is that it would be is smaller.

OpenStudy (anonymous):

*be smaller

OpenStudy (anonymous):

I don't think that's the answer they are looking for though, but thanks

jimthompson5910 (jim_thompson5910):

Let A be the area of the original rectangle So A = LW where L and W are the length and width If the length and width are 1/4 of their original size, then L becomes L/4 and W becomes W/4 So the new area is (L/4)*(W/4) = (LW)/(4*4) = LW/16 = A/16 So the new area is one-sixteenth of the original area.

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

you're welcome